OpenStudy (anonymous):

determine where f is continuous. f(x) = (sin^-1(1/x))/x

OpenStudy (anonymous):

To determind where f(x) is continuous, think about where it is discontinuous. What values is x not allowed to take?

OpenStudy (anonymous):

\[-1 \le x \le 1\]

OpenStudy (anonymous):

understand the answer graphically ...can it be shown algebraically ?

OpenStudy (anonymous):

Yeah you look at the functions involve. Essientially we have\[\sin^{-1} \frac{ 1 }{ x }\] and \[\frac{ 1 }{ x }\] So from the second we get that x cannot be 0. For the first, we know that the domain is -1<=1/x<=1 or -1<=x<=1. Therefore the domain for f(x) is -1<=x<0 and 0<x<=1.

OpenStudy (anonymous):

So basically look at what types of functions are involved, find each of their domains, and then combine them.